OpenStudy (anonymous):

The angle x is the reference angle. We know that cos (3x – 1) = cos (-(3x-1)) = cos (1-3x) I don't get how this works.

OpenStudy (anonymous):

cos (3x – 1) = cos (1-3x) 3x – 1 = 1-3x